A particle is projected vertically up from top of tower with velocity 10m/s. It reaches a ground in 5 sec. Finda. Height of the towerb. Sticking velocity at groundc. Distance traversed by particled. Average speed and average velocity of particle

Grade:11

1 Answers

Jigisha Joshi
6 years ago
distance traversed =85 m/sAvg speed =85/5=17m/sAvg velocity=75/5=15m/sHeight of the tower =S=ut-1/2gt×t=10×5-1/2×10×5×5=50-125=-75mStriking ground velocity=u-gt=10-10×50=-40m/s